# Radio Map Prediction from Aerial Images and Application to Coverage Optimization

### v1

@@ Abstract
In recent years, several studies have explored deep learning algorithms to predict large-scale signal fading, or path loss, in urban communication networks. The goal is to replace costly measurement campaigns, inaccurate statistical models, or computationally expensive ray-tracing simulations with machine learning models that deliver quick and accurate predictions.
We focus on predicting path loss radio maps using convolutional neural networks, leveraging aerial images alone or in combination with supplementary height information. Notably, our approach does not rely on explicit classification of environmental objects, which is often unavailable for most locations worldwide. While the prediction of radio maps using complete 3D environmental data is well-studied, the use of only aerial images remains under-explored. We address this gap by showing that state-of-the-art models developed for existing radio map datasets can be effectively adapted to this task, achieving strong performance. Additionally, we introduce a new model that slightly exceeds the performance of the present state-of-the-art with reduced complexity.
@@ I INTRODUCTION
Important applications of radio maps include the optimization of cellular network layout to achieve good coverage of an area (Section VI), link scheduling [link_scheduling], localization of user equipment (UE) based on RSS measurements at the Rx [locunet] or, in the context of 5G and 6G, beam management [beams]. Since measurement campaigns are impractical and too expensive on a large scale, different methods have been developed to approximate the path loss. Statistical path loss models that express the attenuation solely as a function of the distance between Tx and Rx and additional stochastic effects are inherently unable to capture the radio wave propagation in a specific environment. Ray-tracing simulations provide an option to approximate the underlying physical phenomena with high accuracy [ray-tracing-accuracy]. However, due to the relatively long run times of the simulations and the need for detailed 3D city models, they can be infeasible for applications in real time or on large scale.
@@
This opens up a range of applications in which statistical path loss models are too inaccurate and ray-tracing is too slow. Additionally, DNNs are a very flexible tool for extracting information and patterns implicitly from data. We show that convolutional neural networks (CNN) are capable of implicitly extracting information about the presence, shape, and height of buildings and trees that is necessary to determine radio maps from aerial images.
In brief summary, our main contributions (further discussed in Section II-C) are:
  • We investigate the prediction of radio maps from incomplete environment information in the form of only aerial images or additionally unclassified height data, plus Tx characteristics and map position information via convolutional neural networks (CNN) posed as an image-to-image problem. We propose a novel network architecture based on UNet [unet] and deformable convolutions [dcn] outperforming current state-of-the-art (SOTA) models.
  • We demonstrate through an application example how differentiable radio maps generated by trained CNNs can be used in wireless network optimization.
  • We make a new dataset available that, for the first time, includes consistent aerial images, 3D information of buildings and trees from real places in the world with correct heights and shapes, and ground truth radio maps generated with ray-tracing, featuring directive antennas at the Tx.  [^1] The dataset and some numerical comparisons of model architectures and input features have already been presented in our preprint [jaensch2024radiomapestimation]. For this paper, we have focused more deeply on the investigation of prediction from images, refined our model architecture, and added the investigation of applications to network coverage optimization.
@@
In Section II, we review some necessary wireless communication background to precisely define the investigated problem and give an account of the relevant literature on deep learning for radio map prediction and a comparison to our contributions. Section III describes the generation of the dataset and provides the data sources and simulation parameters we used. Model architectures and the training pipeline are explained in Section IV. In Sections V and VI, we discuss the results on radio map prediction and applications to network coverage optimization, respectively.
II DEFINITIONS, BACKGROUND AND CONTRIBUTION
@@ II-A Wireless Communication Background
Consider a Tx-Rx pair in a fixed environment with $n$ paths establishing the link from the Tx to the Rx. These may include a direct line-of-sight path and multipaths undergoing reflections from surfaces such as the ground or building walls and diffractions around edges or corners of objects. In our simulation scenario described in Section III, we assume that the wavelength of the communication signal is significantly smaller than the accuracy of spatial positions and shapes of objects in the environment. This implies that it is impossible to model the phase differences between different paths appropriately, and any small-scale fading due to constructive and destructive interference between paths arriving with different phase shifts is modeled as a random effect. In fact, in standard wireless communication theory it is customary to separate the large-scale path loss from the small-scale random fading and consider normalized statistics for the latter (e.g., Rayleigh or Rician fading with unit second moment, see [molisch]).
We consider a time-varying uncorrelated scattering model [wssus] with transfer function
@@
$\frac{\operatorname{P}_{\operatorname{R}}}{\operatorname{P}_{\operatorname{T}}}=\sum_{i=1}^{n}\mathbb{E}\left[|c_{p}|^{2}\right],$ (1)
@@
for all $\tau_{p}\neq\tau_{q}$ when $|\tau_{p}-\tau_{q}|>1/W$, i.e., when the frequency variations due to the delay spread are significant over the channel bandwidth. Note also that if two paths have both Doppler shift difference less than $1/T$ and delay difference less than $1/W$, then they can be counted as a single path since they are indistinguishable [proakis]. Hence, the above conditions are always satisfied for all distinguishable paths. We conclude that whether we model the path coefficients as uncorrelated random variables and adopt the definition of average power as an ensemble average, or model them as deterministic quantities and use the definition of average power as a local time-frequency average, the macroscopic path loss is consistently defined as the sum of the squares (or the sum of the second moments) of the path coefficients.
@@
As elaborated in [radiounet], it is reasonable to truncate the path loss from below at a threshold corresponding to the noise floor, as lower power levels are irrelevant in practice, and to work with powers and losses in $\operatorname{dB}$ scale, see Section III-B.
II-B General Approach and Related Works
A common way to model the path loss is to use the log normal shadowing models in different variations (e.g. the 3GPP model described in [3gpp_log_normal]). They originate from the idea to express the path loss merely as a function of the distance between Tx and Rx and to model deviations as the realization of a Gaussian random variable in the log (dB) domain. Albeit the resulting graph of the function together with the stochastic spread may fit measurement points, these models completely disregard the radio wave propagation in a specific environment along specific paths.
Ray-tracing [ray-tracing] provides a more accurate, site-specific solution by modeling electromagnetic waves as rays launched in a very fine subdivision of the space around the Tx. Reflections from surfaces and potentially other effects such as diffractions or transmissions are calculated upon contact with objects [ray-tracing-accuracy]. The rays arriving at the Rx are combined to generate the path loss defined in (1) or other channel state information. As mentioned before, potential drawbacks for certain applications are the runtime of typically at least several seconds to generate a complete radio map and the need to have a complete 3D model of the environment available.
Other methods include completing the radio map from sparse measurements [rm_completion] or estimating the path loss for a single Rx position at a time via ML [ml_single].
In the following, we describe approaches to predict the whole radio map at once using ML. Each pixel in the target radio map, represented by a two-dimensional image, corresponds to a different location in the considered environment from a birds-eye perspective, and its value is the path loss on an appropriate scale. The input information about the presence and potentially height of objects like buildings or trees, the transmitter position, and possibly other parameters that differ from sample to sample are usually encoded in two-dimensional images of the same shape as the target and stacked along an additional channel dimension. All works described in the following build on CNN models [cnn], which have been shown to be effective in other tasks such as semantic segmentation, image denoising, or image-to-image translation that are structurally similar in the sense that input and output of the model are image-like tensors.
A first seminal contribution investigating the prediction of the whole radio map at once using CNNs came out of our research group [radiounet] and considered a UNet [unet] model trained in a supervised manner to approximate radio maps generated with two-dimensional ray-tracing from city maps and Tx locations. The presence or absence of buildings and the Tx are encoded in binary input images. The authors explore strategies to incorporate available signal measurements and adapt the trained models to more realistic scenarios and show applications to coverage estimation and fingerprint localization. Several other works have extended this approach to the more challenging scenario of radio maps generated with 3D ray-tracing, e.g. [plnet, fadenet, radiotrans]. In these works, the binary input information per pixel is usually replaced by height information. Some works have experimented with more variables between the different samples encoded in additional inputs, for example the carrier frequency of the transmitted signal [radiotrans] or antenna patterns and orientations [plnet]. Albeit the basic encoder-decoder with skip-connections design of the UNet has been used in most works since then, it has been soon recognized that standard UNets are inherently limited when it comes to propagating information over longer distances, which is especially important to accurately predict reflections. Several approaches make use of dilated convolutions [dilated] to increase the receptive field [ziemann21, deepray, pmnet], whereas in [radiotrans] vision transformer layers [vit] are added to the CNN to allow modeling long-range relationships. Besides changes to the network architecture, some authors propose feature engineering to resolve the described problem and improve the accuracy. In [qiu22], the network is provided with an input image containing the spatial distance of each pixel position to the transmitter location. The authors of [radiotrans] propose to feed the coordinates of each pixel and the Tx to the network in all positions. To show the validity of training CNNs on simulated data, in [plnet] a model tested first on radio maps generated by ray-tracing is retrained on real-world measurements and shown to perform better than all conventional methods.
Only one work we are aware of investigates path loss prediction for a whole area from aerial images [plgan]. The scenario considered in that study differs from ours as Tx are mounted on unmanned aerial vehicles (UAV) instead of buildings, and the developed model seems rather inaccurate with RMSE over $43\operatorname{dB}$.
The authors of [radiounet] and [pmnet] regard general application of radio maps to coverage area classification and UE localization, which are not specifically requiring DNNs. In [fadenet] it is shown that by making use of the fast inference speed of the trained model, a large number of combinations of potential Tx locations can be evaluated quickly in terms of the coverage provided in an area, reducing the time needed to find the optimal combination significantly compared to using ray-tracing.
Dataset RadioMapSeer [radiounet] RadioMapSeer3D [cagkan3d] USC [pmnet] RMDirectionalBerlin (ours)
Dimension 2D 3D 2D 3D
Antennas (Tx) isotropic isotropic isotropic directional
#Samples $56080$ $50680$ $19016$ $74515$
TABLE I: Overview of open radio map datasets.
II-C Our Contribution
We investigate the prediction of radio maps from aerial images and potentially unclassified height maps for scenarios in which a complete 3D model of the environment is not available. Our results (RMSE of $6.7$dB without height information and $5.2$dB with height information) show a significant improvement over the only other work we are aware of considering a similar scenario ($44$dB in [plgan]).
For this purpose, we have generated city geometries from real places in the city of Berlin and conducted ray-tracing simulations to obtain a large collection of path loss maps modeling typical urban cellular networks, described in Section III. In contrast to the other publicly available radio map datasets we are aware of [radiounet, pmnet, cagkan3d], ours is the first one to feature directional antennas at the Tx and to include trees and realistic building heights with approximated roof shapes taken from the real world as well as aerial images. Table I provides an overview of the similarities and differences between other openly accessible datasets and ours.
We perform experiments with different CNN architectures and input features in order to encode the city geometry, the relative position of the Tx to locations on the map, and the Tx antenna characteristics, including several ideas proposed in the literature (Section IV-B). Lastly, we consider the usage of deformable convolutional layers [dcn] for the radio map prediction task and show that our proposed model achieves the same or higher accuracy as the state of the art with reduced complexity in our scenarios.
In Section VI, we show how a trained model can be used to solve an optimization problem in a cellular network. In contrast to applications in previous works, such as [fadenet], where the fast inference speed of the trained network is used to optimize BS deployment via exhaustive search over all possible combinations of locations, we also take advantage of backpropagation through a trained and frozen network. This allows us to directly optimize input parameters defined as trainable variables with respect to some loss or score function calculated from the predicted radio maps via gradient descent. Although this technique has been used in e.g. inverse material design [nanophotonics_inverse_design], we are not aware of previous usage in our field. While this is also possible with the recently developed ray-tracer Sionna [sionna], our approach does not require a 3D model of the environment and works just with aerial images. Additionally, optimization using our models is significantly faster than with Sionna, which has to recalculate the radio map via ray-tracing in each iteration [sionna_rt].
III DATASET
@@
III-A City Maps
Our goal was to simulate cellular networks in urban environments corresponding to places in the real world and featuring buildings with realistic shapes and heights and also trees. Although a few public institutions provide open 3D models of certain cities, municipalities, or even whole countries, we could not find any that contained information about foliage. However, available height maps or digital surface models (DSMs) lack the necessary classification of objects into buildings, vegetation, and ground. We have therefore decided to generate a new dataset of city maps from raw airborne light detection and ranging (LiDAR) point clouds provided by the Geoportal Berlin [geoportal_berlin]. Using the software LAStools [lastools], we have automatically separated the point clouds into the classes ground, building, and vegetation. The recognition of buildings has been further improved by incorporating building footprints from [geoportal_berlin], categorizing all elevated (above-ground) points inside the footprints as buildings. From there we extracted normalized digital surface models (nDSM), representing the height above the ground for the two classes buildings and vegetation. Since Berlin shows a small deviation in elevation in most areas, we approximated the ground as a flat surface and normalized the heights of objects relative to it.
III-B Radio Map Simulations
The ground truth radio maps were generated with the GPU accelerated X3D propagation model in the widely used ray-tracing software Wireless InSite [wi]. Since it was impossible to determine the exact materials of specific houses or types of trees, we have chosen standard material types for these upon import to the simulation software. The nDSMs have been converted to polygons for the simulations by grouping up neighboring pixels with approximately the same height and interpolating the boundaries. The ground and all buildings are assumed to represent solid structures, allowing reflections from surfaces and diffractions around edges but blocking transmissions. Vegetation on the other hand is modeled as a solely attenuating material. Exploiting that the exact shape of foliage objects is therefore less important, we have chosen less accurate interpolation for the vegetation layers upon constructing the polygons. By doing so, we could significantly reduce the runtime, which depends heavily on the number of faces in the environment geometry, and hence generate a larger dataset.

### v2

@@ Abstract
Several studies have explored deep learning algorithms to predict large-scale signal fading, or path loss, in urban communication networks. The goal is to replace costly measurement campaigns, inaccurate statistical models, or computationally expensive ray-tracing simulations with machine learning models that deliver quick and accurate predictions.
We focus on predicting path loss radio maps using convolutional neural networks, leveraging aerial images alone or in combination with supplementary height information. Notably, our approach does not rely on explicit classification of environmental objects, which is often unavailable for most locations worldwide. While the prediction of radio maps using complete 3D environmental data is well-studied, the use of only aerial images remains under-explored. We address this gap by showing that state-of-the-art models developed for existing radio map datasets can be effectively adapted to this task. Additionally, we introduce a new model dubbed UNetDCN that achieves on par or better performance compared to the state-of-the-art with reduced complexity.
@@ I INTRODUCTION
Important applications of radio maps include the optimization of cellular network layout to achieve good coverage of an area (Section IV), link scheduling [1], localization of user equipment (UE) based on RSS measurements at the Rx [2] or, in the context of 5G and 6G, beam management [3]. Since measurement campaigns are impractical and too expensive on a large scale, different methods have been developed to approximate the path loss. Statistical path loss models that express the attenuation solely as a function of the distance between Tx and Rx and additional stochastic effects are inherently unable to capture the radio wave propagation in a specific environment. Ray-tracing simulations provide an option to approximate the underlying physical phenomena with high accuracy [4]. However, due to the relatively long run times of the simulations and the need for detailed 3D city models, they can be infeasible for applications in real time or on large scale.
@@
This opens up a range of applications in which statistical path loss models are too inaccurate and ray-tracing is too slow, [5], [6]. Additionally, DNNs are a very flexible tool for extracting information and patterns implicitly from data. We show that convolutional neural networks (CNN) are capable of implicitly extracting information about the presence, shape, and height of buildings and trees that is necessary to determine radio maps from aerial images.
@@
In the remainder of this Section, we give an account of the relevant literature for radio map prediction, in particular using deep learning, and a comparison to our contributions. Section II reviews some necessary wireless communication background to precisely define the investigated problem, followed by a description of the generation of the dataset, CNN architectures and other aspects of the experiments. In Section III, we discuss the results on radio map prediction and provide an ablation study for the proposed UNetDCN. Lastly, Section IV showcases applications to network coverage optimization and Section V finishes with the conclusions.
I-A General Approach and Related Works
A common way to model the path loss is to use the log normal shadowing models in different variations (e.g. the 3GPP model described in [7]). They originate from the idea to express the path loss merely as a function of the distance between Tx and Rx and to model deviations as the realization of a Gaussian random variable in the log (dB) domain. Albeit the resulting graph of the function together with the stochastic spread may fit measurement points, these models completely disregard the radio wave propagation in a specific environment along specific paths.
Ray-tracing [8] provides a more accurate, site-specific solution by modeling electromagnetic waves as rays launched in a very fine subdivision of the space around the Tx. Reflections from surfaces and potentially other effects such as diffractions or transmissions are calculated upon contact with objects [4]. The rays arriving at the Rx are combined to generate the path loss as defined in Section II-A, (1), or other channel state information. As mentioned before, potential drawbacks for certain applications are the runtime of typically at least several seconds to generate a complete radio map and the need to have a complete 3D model of the environment available.
Other methods include completing the radio map from sparse measurements [9] or estimating the path loss for a single Rx position at a time via ML [10].
In the following, we describe approaches to predict the whole radio map at once using ML. Each pixel in the target radio map, represented by a two-dimensional image, corresponds to a different location in the considered environment from a birds-eye perspective, and its value is the path loss on an appropriate scale. The input information about the presence and potentially height of objects like buildings or trees, the transmitter position, and possibly other parameters that differ from sample to sample are usually encoded in two-dimensional images of the same shape as the target and stacked along an additional channel dimension. All works described in the following build on CNN models [11], which have been shown to be effective in other tasks such as semantic segmentation, image denoising, or image-to-image translation that are structurally similar in the sense that input and output of the model are image-like tensors.
A first seminal contribution investigating the prediction of the whole radio map at once using CNNs came out of our research group [12] and considered a UNet [13] model trained in a supervised manner to approximate radio maps generated with two-dimensional ray-tracing from city maps and Tx locations. The presence or absence of buildings and the Tx are encoded in binary input images. The authors explore strategies to incorporate available signal measurements and adapt the trained models to more realistic scenarios and show applications to coverage estimation and fingerprint localization. Several other works have extended this approach to the more challenging scenario of radio maps generated with 3D ray-tracing, e.g. [14, 15, 16]. In these works, the binary input information per pixel is usually replaced by height information. Some works have experimented with more variables between the different samples encoded in additional inputs, for example the carrier frequency of the transmitted signal [16] or antenna patterns and orientations [14]. Albeit the basic encoder-decoder with skip-connections design of the UNet has been used in most works since then, it has been soon recognized that standard UNets are inherently limited when it comes to propagating information over longer distances, which is especially important to accurately predict reflections. Several approaches make use of dilated convolutions [17] to increase the receptive field [18, 19, 5], whereas in [16] vision transformer layers [20] are added to the CNN to allow modeling long-range relationships. Besides changes to the network architecture, some authors propose feature engineering to resolve the described problem and improve the accuracy. In [21], the network is provided with an input image containing the spatial distance of each pixel position to the transmitter location. The authors of [16] propose to feed the coordinates of each pixel and the Tx to the network in all positions. To show the validity of training CNNs on simulated data, in [14] a model tested first on radio maps generated by ray-tracing is retrained on real-world measurements and shown to perform better than all conventional methods.
Only one work we are aware of investigates path loss prediction for a whole area from aerial images [22]. The scenario considered in that study differs from ours as Tx are mounted on unmanned aerial vehicles (UAV) instead of buildings, and the developed model seems rather inaccurate with RMSE over $43\operatorname{dB}$.
The authors of [12] and [5] regard general application of radio maps to coverage area classification and UE localization, which are not specifically requiring DNNs. In [15] it is shown that by making use of the fast inference speed of the trained model, a large number of combinations of potential Tx locations can be evaluated quickly in terms of the coverage provided in an area, reducing the time needed to find the optimal combination significantly compared to using ray-tracing. A review on more applications is provided in [6].
Dataset RadioMapSeer [12] RadioMapSeer3D [23] USC [5] RMDirectionalBerlin (ours)
Dimension 2D 3D 2D 3D
Antennas (Tx) isotropic isotropic isotropic directional
#Samples $56080$ $50680$ $19016$ $74515$
TABLE I: Overview of open radio map datasets.
I-B Our Contribution
We investigate the prediction of radio maps from aerial images and potentially unclassified height maps for scenarios in which a complete 3D model of the environment is not available. The radio map is only determined from the information about the environment and the Tx. In particular, contrary to some other works, we do not assume measurements to be available. Our results (RMSE of $6.7$dB without height information and $5.2$dB with height information) show a significant improvement over the only other work we are aware of considering a similar scenario ($44$dB in [22]).
For this purpose, we have generated city geometries from real places in the city of Berlin and conducted ray-tracing simulations to obtain a large collection of path loss maps modeling typical urban cellular networks, described in Section II-B. In contrast to the other publicly available radio map datasets we are aware of [12, 5, 23], ours is the first one to feature directional antennas at the Tx and to include trees and realistic building heights with approximated roof shapes taken from the real world as well as aerial images. Table I provides an overview of the similarities and differences between other openly accessible datasets and ours.  [^1] The dataset and some numerical comparisons of model architectures and input features have already been presented in our preprint [24]. For this paper, we have focused more deeply on the investigation of prediction from images, refined our model architecture, and added the investigation of applications to network coverage optimization. Our dataset can be found at https://zenodo.org/records/13834313, the code at https://github.com/fabja19/RML_v2_img.
We perform experiments with different CNN architectures and input features in order to encode the city geometry, the relative position of the Tx to locations on the map, and the Tx antenna characteristics, including several ideas proposed in the literature (Section II-C2). Lastly, we consider the usage of deformable convolutional layers [25] for the radio map prediction task and show that our proposed model achieves the same or higher accuracy as the state of the art, depending on the scenario, with reduced complexity.
In Section IV, we show how a trained model can be used to solve an optimization problem in a cellular network. In contrast to applications in previous works, such as [15], where the fast inference speed of the trained network is used to optimize BS deployment via exhaustive search over all possible combinations of locations, we also take advantage of backpropagation through a trained and frozen network. This allows us to directly optimize input parameters defined as trainable variables with respect to some loss or score function calculated from the predicted radio maps via gradient descent. Although this technique has been used in e.g. inverse material design [26], we are not aware of previous usage in our field. Note that the optimization of antenna characteristics and materials is also possible with the recently developed ray-tracer Sionna [27]. Our approach, however, does not require a 3D model of the environment and works just with aerial images as inputs, whereas Sionna requires an explicit 3D model as input. Additionally, optimization using our models is significantly faster than with Sionna, which has to recalculate the radio map via ray-tracing in each iteration [28].
II SYSTEM AND METHODOLOGY
@@ II-A Wireless Communication Background
Consider a Tx-Rx pair in a fixed environment with $n$ paths establishing the link from the Tx to the Rx. These may include a direct line-of-sight path and multipaths undergoing reflections from surfaces such as the ground or building walls and diffractions around edges or corners of objects. In our simulation scenario described in Section II-B, we assume that the wavelength of the communication signal is significantly smaller than the accuracy of spatial positions and shapes of objects in the environment. This implies that it is impossible to model the phase differences between different paths appropriately, and any small-scale fading due to constructive and destructive interference between paths arriving with different phase shifts is modeled as a random effect. In fact, in standard wireless communication theory it is customary to separate the large-scale path loss from the small-scale random fading and consider normalized statistics for the latter (e.g., Rayleigh or Rician fading with unit second moment, see [29]).
We consider a time-varying uncorrelated scattering model [30] with transfer function
@@
$\frac{\operatorname{P}_{\operatorname{R}}}{\operatorname{P}_{\operatorname{T}}}=\sum_{p=1}^{n}\mathbb{E}\left[|c_{p}|^{2}\right],$ (1)
@@
for all $\tau_{p}\neq\tau_{q}$ when $|\tau_{p}-\tau_{q}|>1/W$, i.e., when the frequency variations due to the delay spread are significant over the channel bandwidth. Note also that if two paths have both Doppler shift difference less than $1/T$ and delay difference less than $1/W$, then they can be counted as a single path since they are indistinguishable [31]. Hence, the above conditions are always satisfied for all distinguishable paths. We conclude that whether we model the path coefficients as uncorrelated random variables and adopt the definition of average power as an ensemble average, or model them as deterministic quantities and use the definition of average power as a local time-frequency average, the macroscopic path loss is consistently defined as the sum of the squares (or the sum of the second moments) of the path coefficients.
@@
As elaborated in [12], it is reasonable to truncate the path loss from below at a threshold corresponding to the noise floor, as lower power levels are irrelevant in practice, and to work with powers and losses in $\operatorname{dB}$ scale, see Section II-B2.
II-B Dataset
@@
II-B1 City Maps
Our goal was to simulate cellular networks in urban environments corresponding to places in the real world and featuring buildings with realistic shapes and heights and also trees. Although a few public institutions provide open 3D models of certain cities, municipalities, or even whole countries, we could not find any that contained information about foliage. However, available height maps or digital surface models (DSMs) lack the necessary classification of objects into buildings, vegetation, and ground. We have therefore decided to generate a new dataset of city maps from raw airborne light detection and ranging (LiDAR) point clouds provided by the Geoportal Berlin [32]. Using the software LAStools [33], we have automatically separated the point clouds into the classes ground, building, and vegetation. The recognition of buildings has been further improved by incorporating building footprints from [32], categorizing all elevated (above-ground) points inside the footprints as buildings. From there we extracted normalized digital surface models (nDSM), representing the height above the ground for the two classes buildings and vegetation. Since Berlin shows a small deviation in elevation in most areas, we approximated the ground as a flat surface and normalized the heights of objects relative to it.
II-B2 Radio Map Simulations
The ground truth radio maps were generated with the GPU accelerated X3D propagation model in the widely used ray-tracing software Wireless InSite [34]. Since it was impossible to determine the exact materials of specific houses or types of trees, we have chosen standard material types for these upon import to the simulation software. The nDSMs have been converted to polygons for the simulations by grouping up neighboring pixels with approximately the same height and interpolating the boundaries. The ground and all buildings are assumed to represent solid structures, allowing reflections from surfaces and diffractions around edges but blocking transmissions. Vegetation on the other hand is modeled as a solely attenuating material. Exploiting that the exact shape of foliage objects is therefore less important, we have chosen less accurate interpolation for the vegetation layers upon constructing the polygons. By doing so, we could significantly reduce the runtime, which depends heavily on the number of faces in the environment geometry, and hence generate a larger dataset.